OpenStudy (anonymous):

A series circuit contains a 9-volt battery, a 3-ohm resistor and a 2-ohm resistor. What is the voltage drop across the 2-ohm resistor? 3.2 V 3.4 V 3.6 V 3.8 V

OpenStudy (anonymous):

i need help asap

OpenStudy (anonymous):

hello its 3.6 v i=v/r=9/5=1.8 amps then take up the individual resistor and apply v=ir=1.8*2=3.6 over resistor 3, voltage drop is 9-3.6